( xy + y + x + 1 ) dy = dx
Solve this differential equation

Grade:

1 Answers

rahul
9 years ago
dy/dx=1/(x+1)(y+1)
(y+1)dy=dx/x+1
y^2/2+y=ln(x+1)+C